1) In an arithmetic sequence, a4=8 and a7=26. Find a1 and d. Write the first 5 terms of the sequence.

2) find the sum

3) find the sum

4) Find the first

5) Find a7


Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

In an Arithmetic sequence
A[n]= a+(n-1)d
a4=8 =a+3d................................1
a7=26=a+6d..............................2
solve 1 &2
subtract 1 from 2
6d-3d = 18
3d=18
d= 6
a+3d =8
a+3(6)=8
a=-10
First five terms -10,-4,2,8,14



Answer by ikleyn(52778)

.

NEVER post more than one problem in each submission.

It is PROHIBITED by the rules of this forum.